OpenStudy (abbeydavidson22):

A cylindrical container of three rubber balls has a height of 18 centimeters and a diameter of 6 centimeters. Each ball in the container has a radius of 3 centimeters. Find the amount of space in the container that is not occupied by rubber balls. Round your answer to the nearest whole number.

likeabossssssss (likeabossssssss):

welcome to os

OpenStudy (peachpi):

The volume of the cylinder is \(V= \pi r^2h\), where h = 18 cm and r = 3 cm The volume of each ball \(V=\frac{4}{3}\pi r^3\), where r = 3 cm. So find the volume of the cylinder then subtract the volume of three balls to find the volume of the unoccupied space